OpenStudy (anonymous):

When Erik went to bed one winter night in northern Minnesota, the temperature was -2°F. By the time he got up the next morning, the temperature had dropped to -17°F. Assuming the temperature dropped steadily through the night, what was the temperature in the middle of the night? -19°F -9.5°F -15°F -8.5°F

OpenStudy (anonymous):

D, It think!

OpenStudy (hoblos):

the temperature dropped uniformly so at the mid of the night the temperature was at the middle between -2 and -17 which is (-2-17)/2 = -19/2 = -9.5